OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (−6, 3).

OpenStudy (anonymous):

from this perpendicular line u can take the slope of the required line equal to -2/3 then u have a ponit and slope too ... using point slope form u can take the equation of reuired line equation will be like this (y-y1)=m(x-x1) here m is its slope =-2/3 & x1, y1 are the coordinates of the given line i.e. (-6,3)... putting the values of the point coordinates and slope u can take ur line equation